8-5(2)to the 3power/(-8)
Answer by mystic1306(20)

8-5(2)^3
--------
8
first raise ONLY 2 to the 3rd power
2*2*2=8
8-5(8)=
8-40= -32
so you have:
-32
---
-8
divide -32 by -8
remember when you multiply two negitives you get a positive
so -32/-8= 4